# Information theory for data-driven model reduction in physics and biology

### v2

@@ Appendix D Derivations
$p^{*}_{\beta}(h|x)\approx\frac{1}{\mathcal{N}(x)}p^{*}_{\beta}(h)\,\exp\left(\beta\,\text{e}^{\lambda_{1}\Delta t}\phi_{1}(x)f_{1}(h)\right)$ (28)
In Ref. [30], it is shown that the encoding minimizing the IB objective satisfies the implicit equation
$\!\!\!\!p_{\beta}^{*}(h|x)=\frac{p_{\beta}^{*}(h)}{\mathcal{N}(x)}\,\exp\left(-\beta D_{\text{KL}}\left[p_{X_{t+\Delta t}|X_{t}}\big{|}\big{|}p_{X_{t+\Delta t}|H_{t}}\right]\right)$ (29)
where $\mathcal{N}$ is a normalization factor ensuring $\int\text{d}h\,p_{\beta}^{*}(h|x)=1$, and
$p_{\beta}^{*}(h)\equiv p^{*}_{\beta}(h)=\int\text{d}x\,p_{\beta}^{*}(h|x)p(x).$ (30)
@@
$D_{\text{KL}}[p||q]=\int\text{d}x\,p(x)\,\log\frac{p(x)}{q(x)};$ (31)
in the $D_{\text{KL}}$ in (29), the integration is done with respect to the common random variable $X_{t+\Delta t}$ of the two conditional distributions.
Equation 29 contains the conditional distribution $p_{X_{t+\Delta t}|H_{t}}$, which is the composition of two operations: decoding (going from $H_{t}$ to $X_{t}$) and time evolution (going from $X_{t}$ to $X_{t+\Delta t}$). The decoding is performed using Bayes’ theorem $p_{X_{t}|H_{t}}p_{H_{t}}=p_{H_{t}|X_{t}}p_{X_{t}}$, so that we have
$p_{X_{t+\Delta t}|H_{t}}=\int_{X_{t}}p_{X_{t+\Delta t}|X_{t}}\,p_{H_{t}|X_{t}}\,\frac{p_{X_{t}}}{p_{H_{t}}}.$ (32)
Note the appearance of the encoder $p_{H_{t}|X_{t}}$, making (29) only an implicit solution of the optimization problem.
The conditional distribution $p_{X_{t+\Delta t}|X_{t}}$ can be written in terms of right and left eigenvectors of $U$. Here we neglect $U_{\text{ess}}$, which is the operator corresponding to the essential spectrum. The essential spectrum is the part of the spectrum that is not the discrete spectrum; by neglecting it, we are assuming that the essential radius $\rho_{\text{ess}}$ (the maximum absolute value of eigenvalues in the essential spectrum) is small enough compared to the first few eigenvalues $|\lambda_{n}|$ in the point spectrum. While purely deterministic systems may exhibit a large essential radius, the introduction of noise causes the essential spectrum to shrink or disappear [33, 34]. We can then write the conditional distribution in terms of right and left eigenvectors of $U$,
$p_{X_{t+\Delta t}|X_{t}}(x|x^{\prime})$ $=\sum_{n}\text{e}^{\lambda_{n}\Delta t}\rho_{n}(x)\phi_{n}(x^{\prime}).$ (33)
The Kullback-Leibler divergence in (29) can be represented as a sum of two terms,
$D_{\text{KL}}[...]$ $=\int\text{d}x_{t+\Delta t}\,p(x_{t+\Delta t}|x_{t})\,\log p(x_{t+\Delta t}|x_{t})-\int\text{d}x_{t+\Delta t}\,p(x_{t+\Delta t}|x_{t})\,\log p(x_{t+\Delta t}|h_{t}).$ (34)
The first term has no dependence on $h$ and can hence be absorbed into the normalization $\mathcal{N}(x)$. Plugging the decomposition (33) into the second term leads directly to
$p_{\beta}^{*}(h|x)=\frac{1}{\mathcal{N}(x)}\,p_{\beta}^{*}(h)\,\exp\left[\beta\sum_{n}\text{e}^{\lambda_{n}\Delta t}\phi_{n}(x)f_{n}(h)\right]$ (35)
@@
$f_{n}(h)=\int\text{d}x_{t+\Delta t}\,\rho_{n}(x_{t+\Delta t})\log p(x_{t+\Delta t}|h),$ (36)
which can be understood as a quasi-cross entropy between the right eigenvector $\rho_{n}$ and the distribution on $x_{t+\Delta t}$ obtained by decoding $h$. Equation 35 depends only on the existence of the spectral decomposition; we have made no other assumptions on the dynamics (beyond Markovianity) up until this point.
Perturbation theory at the uniform encoder
In this section we derive our main result for the optimal encoder in the limit of high compression (small $\beta$).
Mathematical result – Consider a transfer operator $U$ with infinitesimal generator $\mathcal{L}_{U}$. Assume that $\mathcal{L}_{U}$ has a discrete spectrum with eigenvalues satisfying 0=Re $0=\text{Re }\lambda_{0}>\text{Re }\lambda_{1}>\text{Re }\lambda_{2}\gg\text{Re}\lambda_{3}...$. Then, for $\beta$ just above the first IB transition $\beta_{1}$ so that $f_{n}(h)\rightarrow 0$, the optimal encoder is given approximately by
with corrections due to the second eigenfunction given by $f_{2}(h)\approx f_{1}(h)\text{e}^{-\Gamma\Delta t}+\mathcal{O}(\text{e}^{-2\Gamma\Delta t})$ where $\Gamma=\lambda_{1}-\lambda_{2}>0$ denotes the spectral gap.
@@
$\frac{\partial^{2}}{\partial f_{n}(h_{i})\partial f_{m}(h_{j})}\mathcal{L}_{\text{IB}}.$ (38)
Similar perturbative approaches have been studied in other contexts within IB in [35, 80, 38]. For concreteness, we consider a discrete state $x\in\{1,...,N_{X}\}$ and discrete encoding $h\in\{1,...,N_{H}\}$. We express the encoder as in Eq. 35 where we now expand the normalization terms
$p_{\beta}^{*}(h_{\mu}|x_{i})=\frac{f_{0}(h_{\mu})\exp\left[\beta\sum_{n}\text{e}^{\lambda_{n}\Delta t}\phi_{n}(x_{i})f_{n}(h_{\mu})\right]}{\sum_{h_{\lambda}}f_{0}(h_{\lambda})\exp\left[\beta\sum_{n}\text{e}^{\lambda_{n}\Delta t}\phi_{n}(x_{i})f_{n}(h_{\lambda})\right]}$ (39)
In the following, we simplify notation as
$p(h_{\mu}|x_{i})=\frac{f_{0}^{\mu}\exp\left[\sum_{n}\tilde{\phi}_{n}(x_{i})f_{n}^{\mu}\right]}{\sum_{h_{\lambda}}f_{0}^{\lambda}\exp\left[\sum_{n}\tilde{\phi}_{n}(x_{i})f_{n}^{\lambda}\right]}$ (40)
where $\tilde{\phi}_{n}(x_{i})=\beta\text{e}^{\lambda_{n}\Delta t}\phi_{n}(x_{i})$. For the uniform encoder, $f_{n}^{\mu}=0$ for all $n\neq 0$.
@@
$\partial_{n}^{\mu}p(h_{\nu})$ $=\sum_{x_{i}}p(x_{i})p(h_{\nu}|x_{i})\tilde{\phi}_{n}(x_{i})\left(\delta_{\mu\nu}-p(h_{\mu}|x_{i})\right).$ (41)
To compute the Hessian, we use the fact that $\sum_{\lambda}\partial^{\mu}_{n}p(h_{\lambda})=0$ (by exchanging sum and derivative), and that $p(h_{\mu}|x_{i})|_{\mathbf{f}=0}=p(h_{\mu})$. We note in passing that we find the first derivative terms of $\mathcal{L}_{\text{IB}}$ vanish, see also [36, 38]. The compression term is given by
$\partial_{n}^{\mu}\partial_{m}^{\nu}I(X,H)$ $=\left(\langle\tilde{\phi}_{n}\tilde{\phi}_{m}\rangle-\langle\tilde{\phi}_{n}\rangle\langle\tilde{\phi}_{m}\rangle\right)$
×$\qquad\times\left(\delta_{\mu\nu}p(h_{\mu})-p(h_{\nu})p(h_{\mu})\right)$
@@
$\partial_{n}^{\mu}\partial_{m}^{\nu}I(Y,H)$ $=\left(\left\langle\frac{\langle\tilde{\phi}_{n}\rangle_{p(\cdot,y)}\langle\tilde{\phi}_{m}\rangle_{p(\cdot,y)}}{p(y)^{2}}\right\rangle-\langle\tilde{\phi}_{n}\rangle\langle\tilde{\phi}_{m}\rangle\right)$
×$\qquad\times\left(\delta_{\mu\nu}p(h_{\mu})-p(h_{\nu})p(h_{\mu})\right)$
Angled brackets with no subscript correspond to an average with respect to the steady state distribution, $\langle\cdot\rangle=\sum_{x_{i}}\cdot\,p(x_{i})$. For the second term, we retain the integration variable $y$ for clarity. In sum, the Hessian of the Lagrangian is given by
$H_{(\mu,n),(\nu,m)}$ $=\left(\langle\tilde{\phi}_{n}\tilde{\phi}_{m}\rangle-\beta\left\langle\frac{\langle\tilde{\phi}_{n}\rangle_{p(\cdot,y)}\langle\tilde{\phi}_{m}\rangle_{p(\cdot,y)}}{p(y)^{2}}\right\rangle-(1-\beta)\langle\tilde{\phi}_{n}\rangle\langle\tilde{\phi}_{m}\rangle\right)\left(\delta_{\mu\nu}p(h_{\mu})-p(h_{\nu})p(h_{\mu})\right)$ (42)
which lives in $\mathbb{R}^{(N_{X}\times N_{H})\times(N_{X}\times N_{H})}$; we index $H$ with a multi-index $(\mu,n)$. From the form of Eq. 42, we see that $H$ is given by a Kronecker (tensor) product
@@
Eigenvalues of $H$ are given by products of eigenvalues of $A^{\beta}$ and $G$, while eigenvectors are given by the tensor product of eigenvectors of $A^{\beta}$ and $G$. The appearance of unstable directions of the Hessian corresponds to the appearance of negative eigenvalues in its spectrum. As $G$ does not depend on $\beta$, zero-crossings of eigenvalues of $H$ therefore correspond to zero-crossings of eigenvalues in the spectrum of $A^{\beta}$.
Equilibrium systems: We first study the stability of the matrix $A^{\beta}_{nm}$ in the case of quasi-equilibrium dynamics. In particular, we consider a Fokker-Planck operator $\mathcal{L}_{\text{FP}}$ of the form
$\partial_{t}p(x)=\mathcal{L}_{\text{FP}}p(x)$ $=-\partial_{i}(f_{i}(x)p(x))+D\partial_{i}^{2}p(x).$ (43)
@@
$\mathcal{L}_{\text{FP}}\rho_{0}(x)=-\partial_{i}J_{i}(x)=0$ (44)
@@
×$\quad\quad\quad\quad\times\text{e}^{(\lambda_{j}+\lambda_{m})\Delta t}\langle\phi_{j}\rho_{0},\phi_{m}\rangle$
@@
$A_{nm}=\delta_{nm}\text{e}^{2\lambda_{n}\Delta t}\left(1-\beta\text{e}^{2\lambda_{n}\Delta t}\right)$
@@
$\beta=\text{e}^{-2\lambda_{1}\Delta t},\text{e}^{-2\lambda_{2}\Delta t},...$ (46)
which are increasing in order (remember $\lambda_{i}\leq 0$). It follows that at the first transition, only $f_{1}$ becomes non-zero, and hence the encoder takes the form given by (28). In the equilibrium case, the encoder learns exclusively the first eigenfunction, with no correction due to the second eigenfunction.
General Case: We now show that the first component $f_{1}$ is selected at the first IB transition even when the flux $\mathbf{J}$ is non-zero. From our assumption on the spectrum of $\mathcal{L}_{U}$, namely $0>\text{Re}\lambda_{1}>\text{Re}\lambda_{2}\gg\text{Re}\lambda_{3}...$ it follows that near the first IB transition the IB loss is given by
@@
where we have introduced the shorthand $B_{ij}=\left\langle\frac{\langle\tilde{\phi}_{i}\rangle_{p(\cdot,y)}\langle\tilde{\phi}_{j}\rangle_{p(\cdot,y)}}{p(y)^{2}}\right\rangle$ and all angled brackets denote averaging with respect to the steady state distribution. The stability of the uniform encoder at $f_{n}=0$ is given by the stability of the 2$\times$2 matrix above. The eigenvalues $\omega_{i}$ and eigenvectors $\mathbf{v}_{i}$ of this matrix can be computed explicitly,
@@
$\frac{f_{1}}{f_{2}}$ $=\frac{-\text{e}^{(\lambda_{1}-\lambda_{2})\Delta t}}{2\hat{a}_{12}}\left(-\hat{a}_{11}+\text{e}^{2(\lambda_{2}-\lambda_{1})\Delta t}\hat{a}_{22}-D\right)$ (47)
$=\frac{\text{e}^{\Gamma\Delta t}}{2\hat{a}_{12}}\left(2\hat{a}_{11}+\mathcal{O}(\text{e}^{-2\Gamma\Delta t})\right)$ (48)
$=\text{e}^{\Gamma\Delta t}\frac{\hat{a}_{11}}{\hat{a}_{12}}+\mathcal{O}(\text{e}^{-2\Gamma\Delta t}).$ (49)
@@
$\langle\phi_{n}\phi_{m}\rangle_{\rho_{0}}=-\frac{2}{\lambda_{n}-\lambda_{m}}\langle(\mathbf{J}\cdot\nabla\phi_{n})\phi_{m}\rangle_{\rho_{0}}.$ (50)
@@
Similar to the equilibrium case, we see that at the first transition the encoder depends only on $f_{1}$, giving (28). In contrast to the equilibrium case, there may be a small correction due to the second eigenfunction, however this becomes exponentially small for long times $\Delta t$.
@@
Note that in this calculation, $\phi_{0}(x)$ is constant (following from the assumption of a non-degenerate eigenvalue at 0) so that the $f_{0}(h)$ factors can be absorbed into $p(h)$. This changes in the case of a degenerate ground state, which corresponds a situation where there are decoupled sectors in which the dynamics evolve independently. Then, each eigenfunction corresponding to the zero eigenvalue is piecewise constant on one of the independent sectors. The optimal encoder Eq. 37 will depend instead on $\phi_{0}(x)$, which identifies the independent sectors.

### v3

@@ IV Derivations
$p^{*}_{\beta}(h_{t}|x_{t})\approx\frac{1}{N(x_{t})}p^{*}_{\beta}(h_{t})\,\exp\left(\beta\,\text{e}^{\lambda_{1}\Delta t}\phi_{1}(x_{t})f_{1}(h_{t})\right)$ (S47)
In Ref. [38] (and reproduced above in (S5)), it is shown that the encoding minimizing the IB objective satisfies the implicit equation
$p_{\beta}^{*}(h_{t}|x_{t})=\frac{p_{\beta}^{*}(h_{t})}{N(x_{t})}\,\exp\left(-\beta D_{\text{KL}}\left[p(x_{t+\Delta t}|x_{t})\lVert p(x_{t+\Delta t}|h_{t})\right]\right)$ (S48)
where $N$ is a normalization factor ensuring $\int\text{d}h\,p_{\beta}^{*}(h_{t}|x_{t})=1$, and
$p_{\beta}^{*}(h_{t})\equiv p^{*}_{\beta}(h_{t})=\int\text{d}x_{t}\,p_{\beta}^{*}(h_{t}|x_{t})p(x_{t}).$ (S49)
@@
$D_{\text{KL}}[p||q]=\int\text{d}x\,p(x)\,\log\frac{p(x)}{q(x)}.$ (S50)
In the $D_{\text{KL}}$ in (S48), the integration is done with respect to the common random variable $X_{t+\Delta t}$ of the two conditional distributions.
Equation S48 contains the conditional distribution $p(x_{t+\Delta t}|h_{t})$, which is the composition of two operations: decoding (going from $H_{t}$ to $X_{t}$) and time evolution (going from $X_{t}$ to $X_{t+\Delta t}$). The decoding is performed using Bayes’ theorem $p(x_{t}|h_{t})p(h_{t})=p(h_{t}|x_{t})p(x_{t})$, so that we have
$p(x_{t+\Delta t}|h_{t})=\int\text{d}{x_{t}}p(x_{t+\Delta t}|x_{t})\,p(h_{t}|x_{t})\,\frac{p(x_{t})}{p(h_{t})}.$ (S51)
Note the appearance of the encoder $p(h_{t}|x_{t})$, making (S48) only an implicit solution of the optimization problem.
The conditional distribution $p(x_{t+\Delta t}|x_{t})$ can be written in terms of right and left eigenvectors of $U$. Here we neglect $U_{\text{ess}}$, which is the operator corresponding to the essential spectrum. The essential spectrum is the part of the spectrum that is not the discrete spectrum; by neglecting it, we are assuming that the essential radius $\rho_{\text{ess}}$ (the maximum absolute value of eigenvalues in the essential spectrum) is small enough compared to the first few eigenvalues $|\lambda_{n}|$ in the point spectrum. While purely deterministic systems may exhibit a large essential radius, the introduction of noise causes the essential spectrum to shrink or disappear [49, 50]. We can then write the conditional distribution in terms of right and left eigenvectors of $U$,
$p(x_{t+\Delta t}|x_{t})$ $=\sum_{n}\text{e}^{\lambda_{n}\Delta t}\rho_{n}(x_{t+\Delta t})\phi_{n}(x_{t}).$ (S52)
The Kullback-Leibler divergence in (S48) can be represented as a sum of two terms,
$D_{\text{KL}}[$...] $=\int\text{d}x_{t+\Delta t}\,p(x_{t+\Delta t}|x_{t})\,\log p(x_{t+\Delta t}|x_{t})-\int\text{d}x_{t+\Delta t}\,p(x_{t+\Delta t}|x_{t})\,\log p(x_{t+\Delta t}|h_{t}).$ (S53)
The first term has no dependence on $h_{t}$ and can hence be absorbed into the normalization $N(x)$. Plugging the decomposition (S52) into the second term leads directly to
$p_{\beta}^{*}(h_{t}|x_{t})=\frac{1}{N(x_{t})}\,p_{\beta}^{*}(h_{t})\,\exp\left[\beta\sum_{n}\text{e}^{\lambda_{n}\Delta t}\phi_{n}(x_{t})f_{n}(h_{t})\right]$ (S54)
@@
$f_{n}(h_{t})=\int\text{d}x_{t+\Delta t}\,\rho_{n}(x_{t+\Delta t})\log p(x_{t+\Delta t}|h_{t}),$ (S55)
which can be understood as a quasi-cross entropy between the right eigenvector $\rho_{n}$ and the distribution on $x_{t+\Delta t}$ obtained by decoding $h$. Equation S54 depends only on the existence of the spectral decomposition; we have made no other assumptions on the dynamics (beyond Markovianity) up until this point.
IV.1 Perturbation theory at the uniform encoder
In this section we derive our main result for the optimal encoder in the limit of high compression (small $\beta$). We begin by deriving the result for discrete latent states $h_{t}\in\{0,..,N_{H}-1\}$, before discussing continuous latent variables below. In this section, for notational simplicity we will drop some subscripts and rename $x\equiv x_{t}$, $h\equiv h_{t}$, $y\equiv x_{t+\Delta t}$.
Mathematical result – Consider a transfer operator $U$ with infinitesimal generator $\mathcal{L}_{U}$. Assume that $\mathcal{L}_{U}$ has a discrete spectrum with eigenvalues satisfying $0=\text{Re }\lambda_{0}>\text{Re }\lambda_{1}>\text{Re }\lambda_{2}\gg\text{Re }\lambda_{3}$.... Then, for $\beta$ just above the first IB transition $\beta_{1}$ so that $f_{n}(h)\rightarrow 0$, the optimal encoder is given approximately by
$p^{*}_{\beta}(h|x)=\frac{1}{\mathcal{N}(x)}p^{*}_{\beta}(h)\,\exp\left(\beta\,\text{e}^{\lambda_{1}\Delta t}\phi_{1}(x)f_{1}(h)\right)$ (S56)
with corrections due to the second eigenfunction given by $f_{2}(h)\approx f_{1}(h)\text{e}^{-\Gamma\Delta t}+\mathcal{O}(\text{e}^{-2\Gamma\Delta t})$ where $\Gamma=\lambda_{1}-\lambda_{2}>0$ denotes the spectral gap. For equilibrium systems exhibiting no probability fluxes in the steady state, this result is exact.
@@ IV.1 Perturbation theory at the uniform encoder
$\frac{\partial^{2}}{\partial f_{n}(h_{i})\partial f_{m}(h_{j})}\mathcal{L}_{\text{IB}}.$ (S57)
Similar perturbative approaches have been studied in other contexts within IB in [51, 110, 54]. For concreteness, we consider a discrete encoding $H_{t}\in\{0,...,N_{H}-1\}$. We express the encoder as in Eq. S54 where we now expand the normalization terms
$p_{\beta}^{*}(h_{\mu}|x)=\frac{f_{0}(h_{\mu})\exp\left[\beta\sum_{n}\text{e}^{\lambda_{n}\Delta t}\phi_{n}(x)f_{n}(h_{\mu})\right]}{\sum_{h_{\lambda}}f_{0}(h_{\lambda})\exp\left[\beta\sum_{n}\text{e}^{\lambda_{n}\Delta t}\phi_{n}(x)f_{n}(h_{\lambda})\right]}.$ (S58)
This conditional probability distribution is shorthand for $p_{\beta}^{*}(H_{t}=h_{\mu}|X_{t}=x)$. In the following, we simplify notation as
$p(h_{\mu}|x)=\frac{f_{0}^{\mu}\exp\left[\sum_{n}\tilde{\phi}_{n}(x)f_{n}^{\mu}\right]}{\sum_{h_{\lambda}}f_{0}^{\lambda}\exp\left[\sum_{n}\tilde{\phi}_{n}(x)f_{n}^{\lambda}\right]}$ (S59)
where $\tilde{\phi}_{n}(x)=\beta\text{e}^{\lambda_{n}\Delta t}\phi_{n}(x)$. For the uniform encoder, $f_{n}^{\mu}=0$ for all $n\neq 0$.
@@
$\partial_{n}^{\mu}p(h_{\nu})$ $=\int\text{d}x\,p(x)p(h_{\nu}|x)\tilde{\phi}_{n}(x)\left(\delta_{\mu\nu}-p(h_{\mu}|x)\right).$ (S60)
To compute the Hessian, we use the fact that $\sum_{\lambda}\partial^{\mu}_{n}p(h_{\lambda})=0$ (by exchanging sum and derivative), and that $p(h_{\mu}|x)|_{\mathbf{f}=0}=p(h_{\mu})$. We note in passing that we find the first derivative terms of $\mathcal{L}_{\text{IB}}$ vanish, see also [52, 54]. The compression term is given by
$\partial_{n}^{\mu}\partial_{m}^{\nu}I(X,H)$ $=\left(\langle\tilde{\phi}_{n}\tilde{\phi}_{m}\rangle-\langle\tilde{\phi}_{n}\rangle\langle\tilde{\phi}_{m}\rangle\right)\left(\delta_{\mu\nu}p(h_{\mu})-p(h_{\nu})p(h_{\mu})\right)$
@@
$\partial_{n}^{\mu}\partial_{m}^{\nu}I(Y,H)$ $=\left(\left\langle\langle\tilde{\phi}_{n}\rangle_{p(\cdot|y)}\langle\tilde{\phi}_{m}\rangle_{p(\cdot|y)}\right\rangle_{y}-\langle\tilde{\phi}_{n}\rangle\langle\tilde{\phi}_{m}\rangle\right)\left(\delta_{\mu\nu}p(h_{\mu})-p(h_{\nu})p(h_{\mu})\right)$
Details of the calculations leading to these results are provided in the appendix. Angled brackets with no subscript correspond to an average with respect to the steady state distribution, $\langle\cdot\rangle=\int\text{d}x\,\cdot\,p(x)$. Written out more explicitly, the second term is
$\left\langle\langle\tilde{\phi}_{n}\rangle_{p(\cdot|y)}\langle\tilde{\phi}_{m}\rangle_{p(\cdot|y)}\right\rangle_{y}$ $=\int\text{d}y\,p(y)\left(\int\text{d}x\tilde{\phi}_{n}(x)p(y|x)\right)\left(\int\text{d}x^{\prime}\tilde{\phi}_{m}(x^{\prime})p(y|x^{\prime})\right)$ (S61)
where we retain the integration variable $y$ as a subscript on the angled brackets for clarity. In sum, the Hessian of the Lagrangian is given by
$H_{(\mu,n),(\nu,m)}$ $=\left(\langle\tilde{\phi}_{n}\tilde{\phi}_{m}\rangle-\beta\left\langle\langle\tilde{\phi}_{n}\rangle_{p(\cdot|y)}\langle\tilde{\phi}_{m}\rangle_{p(\cdot|y)}\right\rangle_{y}-(1-\beta)\langle\tilde{\phi}_{n}\rangle\langle\tilde{\phi}_{m}\rangle\right)\left(\delta_{\mu\nu}p(h_{\mu})-p(h_{\nu})p(h_{\mu})\right)$ (S62)
where we index $H$ with a multi-index $(\mu,n)$. From the form of Eq. S62, we see that $H$ is given by a Kronecker (tensor) product
@@
Eigenvalues of $H$ are given by products of eigenvalues of $A^{\beta}$ and $G$, while eigenvectors are given by the tensor product of eigenvectors of $A^{\beta}$ and $G$. Note that $G$ is nothing other than the covariance matrix for a multinomial distribution. The appearance of unstable directions of the Hessian corresponds to the appearance of negative eigenvalues in its spectrum. As $G$ does not depend on $\beta$, zero-crossings of eigenvalues of $H$ therefore correspond to zero-crossings of eigenvalues in the spectrum of $A^{\beta}$.
Equilibrium systems
We first study the stability of the matrix $A^{\beta}_{nm}$ in the case of dynamics which satisfy detailed balance. In particular, we consider a Fokker-Planck operator $\mathcal{L}_{\text{FP}}$ of the form
$\partial_{t}p(x)=\mathcal{L}_{\text{FP}}p(x)$ $=-\partial_{i}(f_{i}(x)p(x))+D\partial_{i}^{2}p(x).$ (S63)
@@
$\mathcal{L}_{\text{FP}}\rho_{0}(x)=-\partial_{i}J_{i}(x)=0$ (S64)
@@
$\quad\quad\quad\quad\times\text{e}^{(\lambda_{j}+\lambda_{m})\Delta t}\langle\phi_{j}\rho_{0},\phi_{m}\rangle$
@@
$A_{nm}=\delta_{nm}\text{e}^{2\lambda_{n}\Delta t}\left(1-\beta\text{e}^{2\lambda_{n}\Delta t}\right)$ (S66)
@@
$\beta=\text{e}^{-2\lambda_{1}\Delta t},\text{e}^{-2\lambda_{2}\Delta t},$...
which are increasing in order (remember $\lambda_{i}\leq 0$). It follows that at the first transition, only $f_{1}$ becomes non-zero, and hence the encoder takes the form given by (S47). In the equilibrium case, the encoder learns exclusively the first eigenfunction, with no correction due to the second eigenfunction.
General Case
We now show that the first component $f_{1}$ is selected at the first IB transition even when the flux $\mathbf{J}$ is non-zero. From our assumption on the spectrum of $\mathcal{L}_{U}$, namely $0>\text{Re}\lambda_{1}>\text{Re}\lambda_{2}\gg\text{Re}\lambda_{3}$... it follows that near the first IB transition the IB loss is given by
@@
where we have introduced the shorthand $B_{ij}=\left\langle\langle\tilde{\phi}_{i}\rangle_{p(\cdot|y)}\langle\tilde{\phi}_{j}\rangle_{p(\cdot|y)}\right\rangle_{y}$ and all angled brackets denote averaging with respect to the steady state distribution. The stability of the uniform encoder at $f_{n}=0$ is given by the stability of the $\times2$ matrix above. The eigenvalues $\omega_{i}$ and eigenvectors $\mathbf{v}_{i}$ of this matrix can be computed explicitly,
@@
$\frac{f_{1}}{f_{2}}$ $=\frac{-\text{e}^{(\lambda_{1}-\lambda_{2})\Delta t}}{2\hat{a}_{12}}\left(-\hat{a}_{11}+\text{e}^{2(\lambda_{2}-\lambda_{1})\Delta t}\hat{a}_{22}-D\right)$
$=\frac{\text{e}^{\Gamma\Delta t}}{2\hat{a}_{12}}\left(2\hat{a}_{11}+\mathcal{O}(\text{e}^{-2\Gamma\Delta t})\right)$
$=\text{e}^{\Gamma\Delta t}\frac{\hat{a}_{11}}{\hat{a}_{12}}+\mathcal{O}(\text{e}^{-2\Gamma\Delta t}).$ (S67)
@@
$\langle\phi_{n}\phi_{m}\rangle_{\rho_{0}}=-\frac{2}{\lambda_{n}-\lambda_{m}}\langle(\mathbf{J}\cdot\nabla\phi_{n})\phi_{m}\rangle_{\rho_{0}}.$ (S68)
@@
Similar to the equilibrium case, we see that at the first transition the encoder depends only on $f_{1}$, giving (S47). In contrast to the equilibrium case, there may be a small correction due to the second eigenfunction, however this becomes exponentially small for long times $\Delta t$.
@@
Note that in this calculation, $\phi_{0}(x)$ is constant (following from the assumption of a non-degenerate eigenvalue at 0) so that the $f_{0}(h)$ factors can be absorbed into $p(h)$. This changes in the case of a degenerate ground state, which corresponds a situation where there are decoupled sectors in which the dynamics evolve independently. Then, each eigenfunction corresponding to the zero eigenvalue is piecewise constant on one of the independent sectors. The optimal encoder Eq. S56 will depend instead on $\phi_{0}(x)$, which identifies the independent sectors.